# Algebraic perturbation theory: traversable wormholes and generalized entropy beyond subleading order

### v1

@@ Abstract
The crossed product has recently emerged as an important tool in high-energy theory. We combine this with another powerful tool, namely pertubation theory, and study the crossed product algebra of a system under a deformation, relating the structure of deformed observables to that of the undeformed theory. In particular, we derive the change in the von Neumann entropy of the type II algebras, and demonstrate that our approach allows one to formally compute this to arbitrarily high orders in perturbation theory. As a concrete example, we apply this machinery to the case of a double-trace deformation of the thermofield double state in AdS/CFT, which is dual to a traversable wormhole in the bulk, obtaining several new contributions to the generalized entropy relative to the original work by Gao, Jafferis, and Wall Gao:2016bin . We comment on the relevance of this framework for black hole evaporation and interiors, as well as on the applicability of the algebraic approach to quantum gravity more generally.
@@ 1 Motivation
$S=\frac{A_{\gamma}}{4G_{\mathrm{N}}}\leavevmode\nobreak\ ,$ (1)
@@
$S_{\mathrm{gen}}=\frac{A}{4G_{\mathrm{N}}}+S_{\mathrm{matter}}\leavevmode\nobreak\ ,$ (2)
@@
However, neither term in (2) is well-defined: in the case of the leading area term, this is due to the need to place a cut-off in the definition of the surface as one approaches the asymptotic boundary, while for the subleading $S_{\mathrm{matter}}$ term, this can be traced to the divergences that arise when attempting to define entanglement entropy in quantum field theory more generally. [^1] See Calabrese:2004eu for a classic review of entanglement entropy in QFT, as well as Rangamani:2016dms ; Headrick:2019eth ; Nishioka:2018khk for reviews which highlight the geometrical relations discussed above, or Hollands:2017dov for a more algebraic treatment. Ultimately, these issues arise from the difficulties of defining entropy in von Neumann algebras of type III, where neither a trace nor reduced density matrices exist. This presents a seemingly fundamental obstacle for QFT, since the algebras of interest are almost always of this type. Phrased more provocatively, the fact that the black hole information paradox exists at all starkly indicates that physicists do not understand entropy as well as we like to think.
@@
In a remarkable recent work Witten:2021unn , Witten proposed a solution to this problem utilizing the crossed product. The basic idea is to consider the type II algebra that results upon adjoining the generator of the modular automorphism group to the original type III algebra Takesaki1973 . Unlike algebras of type III, those of type II do admit a trace, thereby enabling the definition of reduced density matrices and, most interesting for our purposes, a well-defined von Neumann entropy, ${S=-\mathrm{tr}\rho\ln\rho}$ (the relation between the von Neumann and generalized entropies will be discussed later in the text). In light of the fundamental role that entanglement entropy appears to play in black holes and AdS/CFT, it is hoped that a better understanding of this quantity via this more rigorous approach may enable us to make further progress towards quantum gravity. [^2] For example, several works Jefferson:2018ksk ; Jefferson:2019jev ; Raju:2021lwh have pointed out that some of the assumptions underlying the more cavalier approach to entropy generally taken in high-energy theory (e.g., tensor product factorization in defining reduced density matrices) may be problematic, again motivating the need for a different approach.
@@
This crossed product construction has since been generalized to a wide range of systems, including a microcanonical subsector of the TFD Chandrasekaran:2022eqq , de Sitter space Chandrasekaran:2022cip , Rindler space, ball-shaped regions in CFTs, and subregions in AdS/CFT AliAhmad:2023etg , and JT gravity Penington:2023dql ; Kolchmeyer:2023gwa , among others Jensen:2023yxy ; Klinger:2023tgi . Importantly however, as emphasized in AliAhmad:2023etg , this construction relies on fixing the algebra of observables (relative to a given background), meaning that said algebra only contains so-called trivial excitations which do not backreact on the spacetime. However, if we are to understand black hole evaporation, much less quantum gravity more generally, we will certainly need to accommodate changes in the background that shift the defining region of the algebra. As we will show in this work, the crossed product approach naturally handles such deformations of the algebra, enabling us to formally compute the resulting change in observables to arbitrarily high orders in the corresponding perturbation. We regard this as a small but important step towards a perturbative approach to quantum gravity in the algebraic framework. [^3] Of course, there is a rich history within the mathematical physics literature of attempting to synthesize gravity with the operator algebraic approach, however unfortunately aimed primarily at mathematicians; see for example RejznerBook ; Brunetti:2013maa ; Haag:1990ht for complementary works in this vein. In light of the new appreciation for these approaches within high-energy theory, we hope to see greater cross-talk between our two communities in the future.
@@
As an application, we then apply the general perturbative framework we develop to the case of a double-trace deformation of the thermofield double (TFD) state. As shown by Gao, Jafferis, and Wall (GJW) Gao:2016bin , this is dual to a traversable wormhole in the bulk, which causally connects the left and right exteriors of the eternal black hole in AdS for a time set by the strength of the deformation $h$. In that paper, the authors work to linear order in $h$, and argue for a single correction to the subleading $S_{\mathrm{matter}}$ term in the generalized entropy (2). Here, we show how the more rigorous approach from the crossed product subsumes and extends this result. Specifically, we obtain three additional terms that contribute at linear order in $h$, including one at $O(1/N)$ and two at $O(1/N^{2})$, and eight terms that contribute at $O(h^{2})$. Contrary to GJW, we find that two of the terms at linear order in $h$ represent changes to the area term in the generalized entropy (as well as five of the $O(h^{2})$ terms). More abstractly, this demonstrates that the crossed product allows for a controlled treatment of the inherent non-locality of interior operators, in a tractable scenario where everything is in principle computable.
@@
The rest of this paper is organized as follows: in sec. 2, we briefly review how one can define von Neumann entropy via the crossed product construction as presented in Witten:2021unn . Section 3 then introduces our main result, namely a general discussion of how a unitary transformation of the original type III algebra may lead to a change in the type II von Neumann entropy of a subalgebra. At the level of observables, the change takes the form of a weighting of expectation values when computing the trace, which we show can be computed perturbatively to, in principle, arbitrarily high orders in the strength of the perturbation. Then in sec. 4, we apply this general machinery to the example of a double-trace deformation of the traversable wormhole as introduced in GJW Gao:2016bin , and show how our approach allows us to go significantly beyond their results, obtaining eleven new corrections to the generalized entropy up to quadratic order in the strength of the deformation. We close with a discussion of our results in the broader context of quantum gravity and black holes in sec. 5.
@@ 2 Crossed products in AdS-Schwarzschild
$h=h_{r}-h_{l}\leavevmode\nobreak\ .$ (3)
However, the one-sided generators $h_{r,l}$ do not exist as well-defined operators; intuitively, this is because acting with either one alone would generate singular states at the bifurcation surface. For this reason, we will follow AliAhmad:2023etg in referring to the one-sided modular Hamiltonians as modular charges, to avoid confusion with the true modular Hamiltonian (3). Nonetheless, one can effectively adjoin bounded functions of this charge to the algebra $\mathfrak{A}_{r}$ by instead adjoining $T+X$, where $T=h$ is the generator of the canonical outer automorphism group, and $X=h_{l}$ belongs to the commutant. [^5] This still generates singular states in the left algebra, but that is irrelevant for the purposes of the right exterior. Thus, adjoining $T+X$ is formally equivalent to adjoining $h_{r}$. In particular, one constructs the crossed product
$\widehat{\mathfrak{A}}_{r}\coloneqq\mathfrak{A}_{r}\rtimes\mathbb{R}\leavevmode\nobreak\ ,$ (4)
@@
^$\widehat{\mathcal{H}}\coloneqq\mathcal{H}\otimes L^{2}(\mathbb{R})\leavevmode\nobreak\ ,$ (5)
@@
$\widehat{\Psi}\coloneqq\Psi\otimes g(X)^{1/2}\leavevmode\nobreak\ ,$ (6)
with $\Psi$ the vacuum state in $\mathcal{H}$ and $g(X)$ a state in $L^{2}(\mathbb{R})$. For technical reasons, we assume that $g$ is analytic (cf. (33) below). Then for some (but not necessarily all) $\widehat{a}\in\widehat{\mathfrak{A}}$, we may define the trace
$\mathrm{tr}\widehat{a}=\langle{\widehat{\Psi}}|\widehat{a}K^{-1}|{\widehat{\Psi}}\rangle=\int_{-\infty}^{\infty}\!\mathrm{d}X\,e^{X}\left<{\Psi}\right|{\widehat{a}}\left|{\Psi}\right>\leavevmode\nobreak\ .$ (7)
where $K$ is the density state for $\widehat{\mathfrak{A}}$, which Witten:2021unn constructs by decomposing the corresponding modular operator as
$\widehat{\Delta}_{\widehat{\Psi}}=\widetilde{K}K\leavevmode\nobreak\ ,$ (8)
@@
$K=\Delta_{\Psi}e^{-X}g(T+X)=e^{-(T+X)}g(T+X)\qquad\mathrm{and}\qquad\widetilde{K}=e^{X}g(X)^{-1}\leavevmode\nobreak\ ,$ (9)
are the density states on the algebra $\widehat{\mathfrak{A}}$ and its commutant, respectively, and $\Delta_{\Psi}$ is the modular operator on the original type III algebra $\mathfrak{A}$. One can then define the von Neumann entropy on the type II algebra in the usual way, namely
$S(K)=-\mathrm{tr}\,K\ln K\leavevmode\nobreak\ .$ (10)
More generally, one may consider an excited state $\Phi$ obtained by adding matter to the state $\Psi$; the corresponding density state upon performing the crossed product is then Chandrasekaran:2022eqq ,
$K^{\prime}=\epsilon\,\bar{g}\left(\epsilon(T+X)\right)^{1/2}e^{-\beta X}\Delta_{\Phi|\Psi}g\left(\epsilon(T+X)\right)^{1/2}\leavevmode\nobreak\ ,$ (11)
where $\Delta_{\Phi|\Psi}$ is the relative modular operator and $\epsilon$ is a parametrically small constant which controls the semiclassical behavior of the state. Note that when $\Phi=\Psi$, this reduces to ${\Delta_{\Phi|\Psi}=\Delta_{\Psi}}$, and we recover (9). However, a crucial point is that one cannot remain within the same algebra $\widehat{\mathfrak{A}}$ and consider arbitrarily excited states. The basic reason is that fixing an algebra amounts to fixing a region, but an arbitrary excitation may backreact – shifting the region – or lack a geometric interpretation. As explained in AliAhmad:2023etg , the above holds only for so-called trivial excitations which do not affect the geometry of the region. For the more interesting case in which the QES changes under the perturbation, the expectation value becomes weighted relative to the vacuum state; this is the subject of the present work.

### v2

@@ Abstract
The crossed product has recently emerged as an important ingredient in describing algebras of observables for quantum field theory and gravity. We combine this with perturbation theory, and study perturbative crossed product algebras obtained from a unitary deformation of the original system. Motivated by the problem of black hole evaporation, we propose an abstract framework in which black hole information can be transferred to Hawking radiation by passing to a perturbative crossed product exhibiting a degree of non-locality in its modular structure. As both a concrete example and a toy model for evaporation, we analyze the algebra of observables of the traversable wormhole in anti-de-Sitter space. We obtain new contributions to the generalized entropy beyond subleading order relative to the original work by Gao, Jafferis, and Wall Gao:2016bin . We close with some comments on the potential applicability of the algebraic approach to quantum gravity.
@@ 1 Motivation
$S=\frac{A_{\gamma}}{4G_{\mathrm{N}}}\penalty 10000\ ,$ (1)
@@
$S_{\mathrm{gen}}=\frac{A}{4G_{\mathrm{N}}}+S_{\mathrm{matter}}\penalty 10000\ ,$ (2)
@@
However, neither term in (2) is well-defined: in the case of the leading area term, this is due to the need to place a cut-off in the definition of the surface as one approaches the asymptotic boundary, while for the subleading $S_{\mathrm{matter}}$ term, this can be traced to the divergences that arise when attempting to define entanglement entropy in quantum field theory more generally. [^1] See Calabrese:2004eu for a classic review of entanglement entropy in QFT, as well as Rangamani:2016dms ; Headrick:2019eth ; Nishioka:2018khk for reviews which highlight the geometrical relations discussed above, or Hollands:2017dov for a more algebraic treatment. Furthermore, even for black hole horizons – for which $A$ is finite – the denominator of the first term is problematic in the semi-classical limit. Ultimately, these issues arise from the difficulties of defining entropy in von Neumann algebras of type III, where neither a trace nor reduced density operators [^2] In an operator algebra $\mathfrak{A}$, there is a distinction among weights $\omega:\mathfrak{A}\to\mathbb{C}$, density states $\widehat{\omega}$ which are normalizable, vector states $\Psi_{\omega}$ in a Hilbert space, and reduced density matrices or states $\rho_{L}$ which are definable when the Hilbert space $\mathcal{H}$ tensor factorizes as $\mathcal{H}_{\rm L}\otimes\mathcal{H}_{\rm R}$ through a partial trace. However, these mathematical subtleties are not relevant to our main results, so for simplicity we will henceforth use the term “density operator” throughout. exist. This presents a seemingly fundamental obstacle for QFT, since the algebras of interest are almost always of this type.
@@
In a remarkable recent work Witten:2021unn , Witten proposed a solution to this problem utilizing the crossed product. The basic idea is to consider the type II algebra that results upon adjoining the generator of the modular automorphism group to the original type III algebra Takesaki1973 . Unlike algebras of type III, those of type II do admit a trace, thereby enabling the definition of reduced density operators and, most interesting for our purposes, a well-defined von Neumann entropy, $S=-\mathrm{tr}\rho\ln\rho$} (the relation between the von Neumann and generalized entropies will be discussed later in the text). In light of the fundamental role that entanglement entropy appears to play in black holes and AdS/CFT, it is hoped that a better understanding of this quantity via this more rigorous approach may enable us to make further progress towards quantum gravity. [^3] For example, several works Jefferson:2018ksk ; Jefferson:2019jev ; Raju:2021lwh have pointed out that some of the assumptions underlying the more cavalier approach to entropy generally taken in high-energy theory (e.g., tensor product factorization in defining reduced density states) may be problematic, again motivating the need for a different approach. For other work on the relevance of types in AdS/CFT, see e.g. Banerjee:2024fmh ; Banerjee:2023eew .
@@
This crossed product construction has since been generalized to a wide range of systems, including a microcanonical subsector of the TFD Chandrasekaran:2022eqq , de Sitter space Chandrasekaran:2022cip , Rindler space, ball-shaped regions in CFTs, and subregions in AdS/CFT AliAhmad:2023etg , and JT gravity Penington:2023dql ; Kolchmeyer:2023gwa , among others Jensen:2023yxy ; Klinger:2023tgi . Importantly however, as emphasized in AliAhmad:2023etg , this construction relies on fixing the algebra of observables relative to a given background, meaning that said algebra only contains so-called trivial excitations which do not backreact on the spacetime (that is, which do not change the defining region to which the algebra is assigned; for a discussion of diffeomorphism invariance in a related context, see Geng:2024dbl ). However, if we are to understand black hole evaporation, much less quantum gravity more generally, we will certainly need to accommodate changes in the background that shift the defining region of the algebra. As we will show in this work, in the case when the algebras are related by a unitary transformation, the crossed product approach naturally handles such deformations, enabling us to formally compute the deformed entropy to arbitrarily high orders in the corresponding perturbation. We regard this as a small but important step towards a perturbative approach to quantum gravity in the algebraic framework. [^4] Of course, there is a rich history within the mathematical physics literature of attempting to synthesize gravity with the operator algebraic approach; see for example RejznerBook ; Brunetti:2013maa ; Haag:1990ht for complementary works in this vein. In light of the new appreciation for these approaches within high-energy theory, we hope to see greater cross-talk between our two communities in the future.
In particular, we propose a canonical description of quantum black hole evaporation that is entirely formulated in the language of operator algebras. The starting point is the semiclassical algebra of observables for a subregion in a given spacetime, including matter fields and gravitons, without assuming a factorization of the underlying Hilbert space. This accommodates both the gauge–type constraints that obstruct factorization in quantum gravity Casini:2013rba and the fact that a closed gravitational system cannot be split into independent statistical subsystems Raju:2021lwh . The formalism is tailored to produce UV-finite von Neumann entropies by transmuting the type III algebras into their corresponding modular (and thus, semifinite) crossed products. The inputs are:
  1. 1.
    an effective canonical description of matter and propagating gravitons $\mathfrak{A}_{\mathcal{R},0}$ in a subregion $\mathcal{R}$ of a semiclassical spacetime $(M,g)$, consisting of a manifold $M$ and metric $g$, and
  2. 2.
    a dynamical process $U$ that appears non–unitary in the effective description, but which is understood to arise from a fundamentally unitary evolution. This could be a placeholder for the unitary map responsible for black hole evaporation, or some other unitary transformation like an interaction.
Physically, the second ingredient captures semiclassical Hawking radiation in the presence of an evaporating black hole. In a complete theory, this process is expected to be unitary, but in the present approach, the apparent non-unitarity is modeled by the non-locality of the modular Hamiltonian in the effective description of the subsystem on $\mathcal{R}$.
What we mean by this is the following: the local degrees of freedom causally accessible to an observer in $\mathcal{R}$ are those which can be propagated under Hamiltonian evolution and remain in the domain of dependence of $\mathcal{R}$. In AdS/CFT, the spacetime has a conformal boundary which decouples from the effective bulk description. The region $\mathcal{R}$ can be chosen, as an example, to be a spatial slice of $\partial M$ and its domain is the corresponding diamond in the boundary. In the bulk, the local algebra is associated to the region known as the causal wedge Hubeny:2012wa , consisting of causally evolved data from the aforementioned boundary diamond into the bulk. However, it is known that, in AdS/CFT, the boundary subregion $\mathcal{R}$ can reach further than the causal wedge into the entanglement wedge Engelhardt:2014gca ; Dong:2016eik ; Wall:2012uf . In a general background, the latter is properly defined as the domain of support of the modular flow associated to $\mathcal{R}$, and the subalgebra $\mathfrak{A}_{\mathcal{R},0}$ is associated to this entire domain. When the modular and causal flows align, which is very restrictive as in vacuum or symmetric regions, the causal domain and the entanglement wedge are the same. In holography, the entanglement wedge is typically bigger than the causal one, which signals non-locality at the level of the description of local boundary operators on $\mathcal{R}$ only. Of course, evolution on the entire spacetime is assumed to be unitary, but this may appear non-unitary in any effective description of $\mathfrak{A}_{\mathcal{R},0}$.
More specifically, in a general asymptotic description, we may split the modular Hamiltonian of the subregion into a local part corresponding to the simple causal case (say, the one-sided time-translation generator of an eternal black hole) and a non-local part encoding the coupling between the evaporating black hole and the radiation. This non-local part will be treated perturbatively in a deformation parameter $h$ (to be made precise below), while quantum-gravity corrections are organized separately in $\hbar$, $G_{\rm N}$, or $1/N$. The formalism developed here allows one to compute the modular data to any order in these parameters.
@@
As an application, we then apply the general perturbative framework we develop to the case of a double-trace deformation of the thermofield double (TFD) state. As shown by Gao, Jafferis, and Wall (GJW) Gao:2016bin , this is dual to a traversable wormhole in the bulk, which causally connects the left and right exteriors of the eternal black hole in AdS for a time set by the deformation profile. In that paper, the authors work to linear order in $h$, and argue for a single correction to the subleading $S_{\mathrm{matter}}$ term in the generalized entropy (2). Here, we show how the more rigorous approach from the crossed product subsumes and extends this result. Specifically, working up to $O(1/N^{2})$, we obtain five additional terms that contribute at linear order in $h$, and fifteen terms that contribute at $O(h^{2})$. Also contrary to GJW, we find that several of these terms represent changes to the area term in the generalized entropy. More abstractly, this demonstrates that the crossed product allows for a controlled treatment of the inherent non-locality of interior operators, in a tractable scenario where everything is in principle computable.
Before proceeding, let us clarify two potential points of confusion: first, when we apply the crossed product construction to the perturbed algebra, we mean that we start again from the GNS construction on the perturbed state (i.e., one must define a new modular Hamiltonian, for the reasons explained above), not that we apply our results to a time-dependent spacetime. The latter is a more complicated problem, which we comment briefly on in the Discussion; see also Witten:2023xze . We will comment below on how one can cleanly isolate the contributions from the perturbation. A second, related point of confusion stems from the fact that most works cited above apply the crossed product in the specific case where the modular Hamiltonian is the generator of large diffeomorphisms (e.g., the ADM mass, or the Hamiltonian of some idealized observer). In this case, one can interpret the resulting type II algebra as arising from the imposition of gravitational constraints on the original type III theory. However, as demonstrated in AliAhmad:2023etg , the crossed product construction is much more general, and gravity is not intrinsically required to change the type of the algebra; indeed, this has been known to mathematicians since Takesaki’s original work Takesaki1973 . [^5] From different perspectives, the crossed product has been shown to appear in the description of a multitude of systems not manifestly containing dynamical gravity AliAhmad:2024wja ; AliAhmad:2024vdw ; AliAhmad:2024saq ; Klinger:2023tgi ; AliAhmad:2025oli . It is also relevant to point out a significant generalization to locally compact groups $G$ of Takesaki’s theorem on the semifiniteness of the crossed product of a type III factor with the modular automorphism group $\mathbb{R}$ in AliAhmad:2024eun . This is relevant in light of the previous point, because otherwise one might worry that the strength of the perturbations would dominate the graviton fluctuations sometimes interpreted as changing the type. Again however, the latter is more generally due to considering the original type III algebra together with the generator of the corresponding outer automorphism group. Thus, one is free to consider the crossed product applied to the algebra after the deformation, and study its semifinite properties as we do here.
@@
The rest of this paper is organized as follows: in sec. 2, we briefly review how one can define von Neumann entropy via the crossed product construction as presented in Witten:2021unn . Section 3 then introduces our main result, namely a general discussion of the von Neumann entropy of a type II subalgebra following a unitary transformation of the total system spanned by the original type III1 factor and its commutant. At the level of observables, the effect takes the form of a weighting of expectation values when computing the trace, which we show can be computed perturbatively to, in principle, arbitrarily high orders in the strength of the perturbation. Then in sec. 4, we revisit the discussion of black hole evaporation in terms of the framework developed herein. As a concrete example, we apply this general machinery to the example of a double-trace deformation of the traversable wormhole as introduced in GJW Gao:2016bin , and show how our approach allows us to go significantly beyond their results, obtaining twenty new corrections to the entropy up to quadratic order in the strength of the deformation. We close with a discussion of our results in the broader context of quantum gravity and black holes in sec. 5.
@@ 2 Crossed products in AdS-Schwarzschild
$h=h_{r}-h_{l}\penalty 10000\ .$ (3)
However, the one-sided generators $h_{r,l}$ do not exist as well-defined operators; intuitively, this is because acting with either one alone would generate singular states at the bifurcation surface. For this reason, we will follow AliAhmad:2023etg in referring to the one-sided modular Hamiltonians as modular charges, to avoid confusion with the true modular Hamiltonian (3). Nonetheless, one can effectively adjoin bounded functions of this charge to the algebra $\mathfrak{A}_{r}$ by instead adjoining $T+X$, where $T=h$ is the generator of the canonical outer automorphism group, and $X=h_{l}$ belongs to the commutant. [^7] Of course, this still generates singular states in the left algebra, but that is irrelevant for the purposes of the right exterior. Thus, adjoining $T+X$ is formally equivalent to adjoining $h_{r}$. In particular, one constructs the crossed product
$\widehat{\mathfrak{A}}_{r}\coloneqq\mathfrak{A}_{r}\rtimes\mathbb{R}\penalty 10000\ ,$ (4)
@@
$\widehat{\mathcal{H}}\coloneqq\mathcal{H}\otimes L^{2}(\mathbb{R})\penalty 10000\ ,$ (5)
@@
$\widehat{\Psi}\coloneqq\Psi\otimes g(X)^{1/2}\penalty 10000\ ,$ (6)
with $\Psi$ the vacuum state in $\mathcal{H}$ and $g(X)$ a state in $L^{2}(\mathbb{R})$. Then for some (but not necessarily all) $\widehat{a}\in\widehat{\mathfrak{A}}$, we may define the trace [^8] We emphasize that this is only defined up to a state-independent constant, a fact which is crucial in identifying the new algebra as type II; see again Witten:2021unn .
[…]
where $K$ is the density operator for $\widehat{\mathfrak{A}}$ in the defining state, which Witten:2021unn constructs by decomposing the corresponding modular operator as
$\widehat{\Delta}_{\widehat{\Psi}}=\widetilde{K}K\penalty 10000\ ,$ (8)
@@
$K=\Delta_{\Psi}e^{-X}g(T+X)=e^{-(T+X)}g(T+X)\qquad\mathrm{and}\qquad\widetilde{K}=e^{X}g(X)^{-1}\penalty 10000\ ,$ (9)
are the density operators on the algebra $\widehat{\mathfrak{A}}$ and its commutant, respectively, and $\Delta_{\Psi}$ is the modular operator on the original type III algebra $\mathfrak{A}$. One can then define the von Neumann entropy on the type II algebra in the usual way, namely
$S(K)=-\mathrm{tr}\,K\ln K\penalty 10000\ .$ (10)
More generally, one may consider an excited state $\Phi$ obtained by adding matter to the state $\Psi$; the corresponding density operator upon performing the crossed product is then Chandrasekaran:2022eqq ,
$K^{\prime}=\epsilon\,\bar{g}\left(\epsilon(T+X)\right)^{1/2}e^{-\beta X}\Delta_{\Phi|\Psi}g\left(\epsilon(T+X)\right)^{1/2}\penalty 10000\ ,$ (11)
where $\Delta_{\Phi|\Psi}$ is the relative modular operator and $\epsilon$ is a parametrically small constant which controls the semiclassical behavior of the state; see Penington:2023dql ; Kudler-Flam:2023hkl for further discussion. Note that when $\Phi=\Psi$, this reduces to $\Delta_{\Phi|\Psi}=\Delta_{\Psi}$}, and we recover (9). However, a crucial point is that one cannot consider arbitrarily excited states while remaining within the same algebra $\widehat{\mathfrak{A}}$. The basic reason is that fixing an algebra amounts to fixing a region, but an arbitrary excitation may backreact – shifting the region – or lack a geometric interpretation. Thus, as explained in AliAhmad:2023etg , the above holds only for so-called trivial excitations which do not backreact sufficiently to affect the geometry of the region. For the more interesting case in which the QES demarcating the region changes under the perturbation, the expectation value becomes weighted relative to the vacuum state; this is the subject of the present work.